Grade 12Physical Chemistry

relation between inversion temperature, critical temperature and boyle temperature

Aman
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student,

µJT = ∂T/∂p
At the inversion temperature the coefficient changes its sign.
For higher temperatures it is negative and the gas heats up due to expansion. For a temperature below the inversion temperature µJT is positive and the expanding gas cools down.
The importance of this temperature is in gas liquefaction. Processes like the Linde process for air liquefaction need a positive coefficient. Hence you can only operate at temperatures below the inversion temperature.

Relation between inverion temperature and Boyle`s Temperature is as follows:Ti is 2a/Kb and Tb is a/Kb so Ti/Tb=2/1. Hence Ti =2Tb .